OpenStudy (anonymous):

After the fraction x plus 2 all over 3 minus the fraction x plus 1 all over 2x have been combined using the least common denominator of 6x, what is the numerator? @SolomonZelman @midhun.madhu1987 Help Please

OpenStudy (solomonzelman):

Can you draw the question please ?

OpenStudy (anonymous):

Alright is going to take some time

OpenStudy (midhun.madhu1987):

draw???

OpenStudy (anonymous):

\[\frac{ x+2 }{ 3 }-\frac{ x+1 }{ 2x}\]

OpenStudy (solomonzelman):

this is the problem ? you are just subtracting fractions ?

OpenStudy (anonymous):

Read the question Again

OpenStudy (anonymous):

Yes that is the problem and you are subtracting

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{(x+2)}{3} - \frac{(x+1)}{2x} }\) \(\LARGE\color{black}{ \frac{(x+2)\color{red}{\times2x}}{3\color{red}{\times2x}} - \frac{(x+1)\color{red}{\times 3}}{2x \color{red}{\times3}} }\) \(\LARGE\color{black}{ \frac{2x(x+2)}{6x} - \frac{3(x+1)}{6x} }\) \(\LARGE\color{black}{ \frac{2x(x+2)-3(x+1)}{6x} }\) \(\LARGE\color{black}{ \frac{2x^2+4x-3x-3}{6x} }\) \(\LARGE\color{black}{ \frac{2x^2+x-3}{6x} }\)

OpenStudy (anonymous):

Good thanks can help with: What is the least common denominator of the fraction 7 over 3 x plus the fraction 2 over x plus the fraction x over x minus 1 ?

OpenStudy (anonymous):

Can you help